OpenStudy (anonymous):

Simplify the given expression. square root of negative 108 end root plus square root of negative 27

OpenStudy (anonymous):

my choices are negative 3 square root of 15 3 I square root of 15 9 I square root of 3 negative 9 square root of 3

OpenStudy (anonymous):

This is the expression we're told to simplify: \[\sqrt{-108}+\sqrt{-27}\] Whenever we have a negative value under a radical sign, we need to pull out a -1, whose square root is i. We then have: \[i \sqrt{108} + i \sqrt{27}\] From the 108, we can pull out the perfect square 36 (whose square root is 6) and from the 27 we can pull out the perfect square 9 (whose square root is 3) so we can simplify this to: \[6i \sqrt{3} + 3i \sqrt{3}\] Combine like terms and we get our answer: \[9i \sqrt{3} \]